The Future of Scientific Computing for Grand Challenge Problems

Computation Structures Group Memo 383 June 11, 1996

Arvind

Andrew Shaw

John Marshall

Submitted to Science

This report describes research done at the Laboratory for Computer Science and the Department of Earth, Atmospheric, and Planetary Sciences of the Massachusetts Institute of Technology. Funding for the Laboratory for Computer Science is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-92-J-1310.

1